OpenStudy (anonymous):

Suppose the velocity of money is constant. The level of output in the economy is Y = 1000 and the price level is equal to CPI = 100. Suppose the money supply changes from M = 2 to M = 4. According to our understanding of the quantity theory of money which is the most likely outcome? A. Y changes to 500 B. Y changes to 2000 C. CPI changes to 200 D. CPI changes to 50 E. The CPI changes to 200 and Y changes to 2000

OpenStudy (anonymous):

mv=py if v is constant, then for a particular m, we have a negative slope graph for p vs y as m changes, the graph moves towards right..hence the value for both y and p changes to retain equilibrium..hence (d)

OpenStudy (anonymous):

The relationship between velocity, the money supply, the price level, and output is represented by the equation M * V = P * Y where M is the money supply, V is the velocity, P is the price level, and Y is the quantity of output. P * Y, the price level multiplied by the quantity of output, gives the nominal GDP. This equation can thus be rearranged as V = (nominal GDP) / M. Conceptually, this equation means that for a given level of nominal GDP, a smaller money supply will result in money needing to change hands more quickly to facilitate the total purchases, which causes increased velocity. The equation for the velocity of money, while useful in its original form, can be converted to a percentage change formula for easier calculations. In this case, the equation becomes (percent change in the money supply) + (percent change in velocity) = (percent change in the price level) + (percent change in output). The percentage change formula aids calculations that involve this equation by ensuring that all variables are in common units. In ur example the Velocity of money remains constant then percent change in velocity = 0% and the money supply has been twiced it means that it has increased by 100%. Therefore, (% change in the money supply) + (% change in velocity) = (% change in the price level) + (% change in output) 100 % + 0% = (% change in the price level) + (% change in output) it means that RHS should be equal to 100 And as we know that if the money inflow into the market increases the price of the commodity increases hence the output

OpenStudy (anonymous):

thus to achieve equilibrium the Y changes to 2000 ( increases by 100%)